OpenStudy (anonymous):

x^2-5x+6/x-2 identify the holes

terenzreignz (terenzreignz):

Now, the first thing you need to do is factor the numerator. Just to see stuff more clearly...

terenzreignz (terenzreignz):

I mean, clearly, the denominator can no longer be factored...

OpenStudy (anonymous):

could you like step by step show me how to solve this please

terenzreignz (terenzreignz):

I'm not really good at teaching how to factor... suffice it to say, the numerator factored is (x - 2)(x - 3)

terenzreignz (terenzreignz):

\[\Large =\frac{(x-2)(x-3)}{x-2}\] right?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

Now, a 'hole' is a point where BOTH the numerator and denominator are zero. So... for what value(s) of x is the *denominator* equal to zero?

OpenStudy (anonymous):

im not sure is it?

terenzreignz (terenzreignz):

What is the denominator?

OpenStudy (anonymous):

x-2

terenzreignz (terenzreignz):

And when is this equal to zero? when x = ...?

OpenStudy (anonymous):

-2

terenzreignz (terenzreignz):

When x = -2 then... x - 2 becomes -2 - 2 = -4 which isn't zero :P ...try again

OpenStudy (anonymous):

-3 as one of them

terenzreignz (terenzreignz):

No... lol For x - 2 = 0 Just solve for x...

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

Yes, 2. So the denominator becomes zero when x = 2 Now what about the numerator? When is the numerator (x-2)(x-3) = 0?

OpenStudy (anonymous):

2,3

terenzreignz (terenzreignz):

That's right. So, numerator is zero when x is 2 or 3 denominator is zero when x is 2... so.... when are they BOTH zero?

OpenStudy (anonymous):

when x is 2 or 3?

terenzreignz (terenzreignz):

No... because when x = 3, the denominator is not zero :P

OpenStudy (anonymous):

=/

OpenStudy (anonymous):

2,3 is wrong

terenzreignz (terenzreignz):

Yup. I told you, holes would be the values of x that would make BOTH the numerator and denominator zero.

OpenStudy (anonymous):

ok so what do i have to do

terenzreignz (terenzreignz):

Think of it this way... the set of values for x that would make the numerator zero is {2,3} while the set of x-values that make the denominator zero is {2} The elements they have in common, that's the answer.